OpenStudy (anonymous):

Solve x/x+4-4/-x=x^2+16/x^2-16

OpenStudy (anonymous):

is this x/(x+4) - 4/(-x) = x^2 + 16(x^2-16) ?

myininaya (myininaya):

\[\frac{x}{x+4}-\frac{4}{-x}=\frac{x^2+16}{x^2-16}\] ?

myininaya (myininaya):

did i right correctly or did i right it wrong?